# NNN: Next-Generation Neural Networks for Marketing Mix Modeling

### v2

@@ 13.1 Future Directions
The promising results of this initial exploration of NNN suggest several important avenues for future research and development aimed at extending its capabilities and robustness:
Incorporating Cost Data for ROAS and Optimization. The current work focused on attribution based on media volume (e.g., impressions). A critical next step is to integrate media spend data, enabling the direct calculation of Return on Ad Spend (ROAS). Furthermore, exploring methods to apply marketing budget optimization techniques based on NNN’s outputs would significantly enhance its practical applicability for marketers.
@@
Validation through Live Experimentation. While simulations and retrospective analyses are valuable, the practical utility of NNN must be confirmed through rigorous, live marketing experiments. Comparing NNN’s performance, attribution results, and potential optimization recommendations directly against established traditional MMMs in real-world settings combined with live experimentation is essential for comprehensive validation and identifying areas requiring further improvement.
@@ 14 Conclusion
This paper introduced NNN, a novel framework for Marketing Mix Modeling (MMM) leveraging the capabilities of neural networks, specifically Transformer architectures, and rich embedding representations for both paid media channels and organic signals like Google Search. Our approach utilizes the flexibility of neural networks to capture complex interactions and temporal dynamics, employs embeddings to integrate quantitative and qualitative channel information, and adopts common machine learning practices, such as optimization via L1 regularization and model selection based on out-of-distribution (OOD) predictive performance.
Through evaluations on both simulated and real-world datasets, we demonstrated NNN’s potential. Our results indicate considerable improvements in predictive accuracy compared to traditional MMM benchmarks, particularly when qualitative data aspects represented by embeddings are significant. Furthermore, NNN yielded plausible sales attribution estimates, on par with traditional MMM. Key aspects of our framework, such as the explicit modeling of intermediary channels like organic Search, offer a mechanism to capture and attribute marketing effects potentially missed by conventional methods. Moreover, the trained NNN model naturally lends itself as a flexible platform for diverse simulation-based analyses; for instance, evaluating the inferred effectiveness of specific creatives or keywords (i.e., creative analysis) can be performed directly within the framework, complementing traditional A/B testing. Finally, the model selection process, focused on tuning a single regularization parameter based on OOD performance, presents a potentially more streamlined workflow compared to iterative prior adjustments common in Bayesian MMMs.
While the results are promising, NNN is a novel methodology and should be viewed in context. It is not yet a direct replacement for established traditional MMM frameworks, which benefit from decades of practical application, refinement, and established best practices. We hope our methods and results inspire applications of neural networks in media mix modeling, live experiments of these techniques, and continued research in this area.
@@ Acknowledgements
Thank you to Karima Zmerli for your steadfast support of this research, and to Jennifer Steele for the early conversations about using Search as a proxy for consumer demand. We would like to thank everyone in the Google Data Science community not explicitly mentioned above for their continued support and guidance.

### v3

@@ 13.1 Future Directions
The promising results of this initial exploration of NNN suggest several important avenues for future research and development aimed at validating the methodology and extending its capabilities and robustness:
Additional Validation Required While simulations and retrospective analyses are valuable, the practical utility of NNN must be confirmed through rigorous, live marketing experiments. Comparing NNN’s performance, attribution results, and potential optimization recommendations directly against established MMMs in real-world settings combined with live experimentation is essential for comprehensive validation and identifying areas requiring further improvement. While this gives us promising hints, but significant additional research and validation is required (especially on real-world datasets) in order to fully understand the tradeoffs and potential risks of this framework.
Incorporating Cost Data for ROAS, and Optimization. The current work focused on attribution based on media volume (e.g., impressions). A critical next step is to integrate media spend data, enabling the direct calculation of Return on Ad Spend (ROAS). Furthermore, exploring methods to apply marketing budget optimization techniques based on NNN’s outputs would significantly enhance its practical applicability for marketers.
@@
Incorporating channels without embeddings Currently NNN has only been tested with a two-channel framework (Search Ads and YouTube) using embedding data. This contrasts with MMMs which can model the effectiveness of dozens of channels at once. Further research is required to understand to what extent NNN can handle additional channels, especially channels that lack embedding data.
@@ 14 Conclusion
This paper introduced NNN, a novel framework for Marketing Measurement leveraging the capabilities of neural networks, specifically Transformer architectures, and rich embedding representations for both paid media channels and organic signals like Google Search. Our approach utilizes the flexibility of neural networks to capture complex interactions and temporal dynamics, employs embeddings to integrate quantitative and qualitative channel information, and adopts common machine learning practices, such as optimization via L1 regularization and model selection based on out-of-distribution (OOD) predictive performance.
Through evaluations on both simulated and real-world datasets, we demonstrated NNN’s potential. Our initial examples yielded plausible sales attribution estimates, on par with a baseline MMM. Additionally, when using NNN’s framework of OOD performance for model assessment, we see considerable improvements in predictive accuracy relative to a baseline MMM. Key aspects of our framework, such as the explicit modeling of intermediary channels like organic Search, offer an exciting potential mechanism to capture and attribute marketing effects potentially missed by conventional methods. Moreover, the trained NNN model naturally lends itself as a flexible platform for diverse simulation-based analyses; for instance, evaluating the inferred effectiveness of specific creatives or keywords (i.e., creative analysis) can be performed directly within the framework, complementing traditional A/B testing. Finally, the model selection process, focused on tuning a single regularization parameter based on OOD performance, presents a potentially streamlined workflow for model development and selection.
While the results are promising, NNN is a novel research direction and should be viewed in context. It is not a direct replacement for currently established MMM frameworks, which benefit from decades of practical application, refinement, and established best practices. We see NNN as a complement rather than a replacement to MMM, with additional capabilities for channels with embedding data but a greater need for research and understanding of its uncertainties. We hope our methods and results inspire applications of neural networks in media measurement, live experiments of these techniques, and continued research in this area.
@@ Acknowledgements
Thank you to Nithya Mahadevan and Karima Zmerli for your steadfast support of this research and for invaluable advice along the way, and to Jennifer Steele for the early conversations about using Search as a proxy for consumer demand. We would like to thank everyone in the Google Data Science community not explicitly mentioned above for their continued support and guidance.